The tape in a videotape cassette has a total length 241 m and can play for 2.1 h. As the tape starts to play, the full reel has an outer radius of 49 mm and an inner radius of 14 mm At some point during the play, both reels will have the same angular speed What is this common angular speed? Answer in units of rad/s.
Answer #1

Solution)

Here, 2(Rf^2 - r^2) = R^2 - r^2

Rf= 0.036 m

Now, linear speed v

v = 241/2.1m/h = 114.7 m/h =0.0318 m/s

Now, angular speed

w = 0.0318/0.036 rad/s = 0.885 rad/s (Ans)

======

Good luck!:)